STA2023-TIR 12:30-t45 (15859) 5Spg 2019 192) Time Romaining 11326 An existing inventory for a test measuring self-esteem indicates that the scores have a standard deviation of 10. A psychologist gave the sef-esteem test to a of 70 individuals, and their mean score was 65. Construct a 95% confidence itenal for the true mean or ai test scoes. table below, to at least three decimal places. Round your answers to one decimal place (If necessary, consult a list of formulas.) what is the lower limit of the 95% confidence interval? what is the upper limit of the 95% confdence interval? U Worked Example 9 6 3
Answer #1

Solution:-

Lower limit = 62.7

Upper limit = 67.3

95% confidence interval for the mean is C.I = ( 62.7, 67.3).

C.I = \bar{x}\pm z_{\alpha /2}\ast \frac{\sigma }{\sqrt{n}}

C.I = 65 + 1.96*1.19523

C.I = 65 + 2.3426

C.I = ( 62.657, 67.343)